A student exerts a force of 140 N at an angle of 30 above the horizontal on a box and causes it to move 2.5 m horizontally.what is the work done on the box?


1
Expert's answer
2020-03-30T08:05:41-0400
"W=Fs\cos{30}"

"W=(140)(2.5)\cos{30}=303\ J"